### v1

@@ Abstract
Index Terms:
Wi-Fi, Channel quality prediction, Machine Learning, Convolutional Neural Networks (CNN), Recurrent Neural Networks, Long Short-Term Memory (LSTM), Bidirectional LSTM (Bi-LSTM).
@@ IV-C Selected hyperparameters
Ch. Ch. $[\cdot 10^{-3}]$ [%] [%]
$1$ $1$ LSTM 2.08 5.59 11.68 30.00 99.90 2.77 3.67 7.48 10.82 17.35 31.50 -31.60 -6.39 4.67 31.50
$1$ All LSTM 1.98 5.45 11.50 29.60 99.50 2.68 3.60 7.42 10.70 17.20 31.40 -31.50 -6.30 4.60 31.30
$5$ $5$ LSTM 1.15 1.57 2.69 17.50 176.30 2.16 2.63 3.95 5.18 13.23 41.95 -42.10 -5.21 3.60 42.10
$5$ All LSTM 1.10 1.51 2.63 17.35 175.75 2.10 2.61 3.90 5.12 13.12 41.80 -41.85 -5.15 3.55 41.70
$9$ $9$ LSTM 4.21 5.91 7.80 60.40 247.90 4.88 4.27 7.68 8.85 24.60 49.90 -49.95 -2.12 6.00 49.95
$9$ All LSTM 4.12 5.85 7.72 60.18 247.40 4.82 4.24 7.62 8.80 24.50 49.80 -49.88 -2.08 5.96 49.85
@@
Ch. Ch. $[\cdot 10^{-3}]$ [%] [%]
$1$ $1$ Bi-LSTM 2.07 5.57 11.65 29.95 99.85 2.76 3.66 7.46 10.80 17.30 31.45 -31.55 -6.37 4.65 31.45
$1$ All Bi-LSTM 1.97 5.50 11.48 29.50 99.45 2.70 3.60 7.42 10.70 17.18 31.35 -31.50 -6.32 4.62 31.32
$5$ $5$ Bi-LSTM 1.14 1.56 2.68 17.48 176.00 2.14 2.62 3.94 5.15 13.20 41.80 -41.85 -5.18 3.58 41.85
$5$ All Bi-LSTM 1.11 1.53 2.65 17.32 175.70 2.12 2.60 3.92 5.12 13.15 41.75 -41.80 -5.15 3.55 41.70
$9$ $9$ Bi-LSTM 4.19 5.90 7.78 60.38 247.50 4.87 4.26 7.67 8.83 24.58 49.88 -49.90 -2.11 5.98 49.88
$9$ All Bi-LSTM 4.08 5.84 7.70 60.12 247.20 4.80 4.23 7.63 8.78 24.50 49.75 -49.85 -2.08 5.95 49.82

### v2

@@ IV-C Selected hyperparameters
Ch. Ch. $\cdot 10^{-3}] [$%] [%]
11 11 LSTM 2.08 5.59 11.68 30.00 99.90 2.77 3.67 7.48 10.82 17.35 31.50 -31.60 -6.39 4.67 31.50
11 All LSTM 1.98 5.45 11.50 29.60 99.50 2.68 3.60 7.42 10.70 17.20 31.40 -31.50 -6.30 4.60 31.30
55 55 LSTM 1.15 1.57 2.69 17.50 176.30 2.16 2.63 3.95 5.18 13.23 41.95 -42.10 -5.21 3.60 42.10
55 All LSTM 1.10 1.51 2.63 17.35 175.75 2.10 2.61 3.90 5.12 13.12 41.80 -41.85 -5.15 3.55 41.70
99 99 LSTM 4.21 5.91 7.80 60.40 247.90 4.88 4.27 7.68 8.85 24.60 49.90 -49.95 -2.12 6.00 49.95
99 All LSTM 4.12 5.85 7.72 60.18 247.40 4.82 4.24 7.62 8.80 24.50 49.80 -49.88 -2.08 5.96 49.85
@@
Ch. Ch. $\cdot 10^{-3}] [$%] [%]
11 11 Bi-LSTM 2.07 5.57 11.65 29.95 99.85 2.76 3.66 7.46 10.80 17.30 31.45 -31.55 -6.37 4.65 31.45
11 All Bi-LSTM 1.97 5.50 11.48 29.50 99.45 2.70 3.60 7.42 10.70 17.18 31.35 -31.50 -6.32 4.62 31.32
55 55 Bi-LSTM 1.14 1.56 2.68 17.48 176.00 2.14 2.62 3.94 5.15 13.20 41.80 -41.85 -5.18 3.58 41.85
55 All Bi-LSTM 1.11 1.53 2.65 17.32 175.70 2.12 2.60 3.92 5.12 13.15 41.75 -41.80 -5.15 3.55 41.70
99 99 Bi-LSTM 4.19 5.90 7.78 60.38 247.50 4.87 4.26 7.67 8.83 24.58 49.88 -49.90 -2.11 5.98 49.88
99 All Bi-LSTM 4.08 5.84 7.70 60.12 247.20 4.80 4.23 7.63 8.78 24.50 49.75 -49.85 -2.08 5.95 49.82